# Differentiability of solutions for a degenerate fully nonlinear free transmission problem

### v1

@@ 2.2 Elementary notions and auxiliary results
3 The existence of solutions
The study of the existence of solutions for fully nonlinear free transmission problems entails genuine difficulties. For instance, fundamental properties - such as properness - are not available for the operator. As a consequence, the comparison principle for viscosity solutions and Perron’s method are out of reach. To establish the existence of a viscosity solution to (1) equipped with a Dirichlet boundary condition, we resort to the approximation strategy devised in [44, 25]. In the sequel, we introduce an approximate equation.
Let $v\in C(\overline{\Omega})$. For $\varepsilon\in(0,1)$, define $g_{v,\varepsilon}:\mathbb{R}^{d}\to[0,1]$ as
$g_{v,\varepsilon}:=\begin{cases}\max\left(\min\left(\frac{v+\varepsilon}{2\varepsilon},1\right),0\right)&\text{ in }\Omega,\\ 0&\text{ in }\mathbb{R}^{d}\setminus\Omega.\end{cases}$
The heuristic underlying the definition of $g_{v,\varepsilon}$ is the following. In $\{v>\varepsilon\}$, we have $g_{v,\varepsilon}=1$, whereas $g_{v,\varepsilon}=0$ in the set $\{v<-\varepsilon\}$. In the region $\{-\varepsilon\leq v\leq\varepsilon\}$ the function takes values between 00 and $1$. The idea behind $g_{v,\varepsilon}$ is to build an indicator function outside of an $\varepsilon$-neighborhood of the free boundary. To smoothen $g_{v,\varepsilon}$, we define a weight function $h_{v,\varepsilon}$ as follows: for $x\in\Omega$, set
$h_{v,\varepsilon}(x):=\left(g_{v,\varepsilon}*\eta_{\varepsilon}\right)(x).$
Then we introduce the approximating degeneracy function $\sigma_{v,\varepsilon}$, given by
$\sigma_{v,\varepsilon}(p,x):=\sigma_{1}(p)h_{v,\varepsilon}(x)+(1-h_{v,\varepsilon}(x))\sigma_{2}(p).$
Remark 1.
Because $0\leq g_{v,\varepsilon}\leq 1$, we have $0\leq h_{v,\varepsilon}\leq 1$. Therefore, $\sigma_{v,\varepsilon}$ is a convex combination of $\sigma_{1}$ and $\sigma_{2}$. Hence, we conclude $\sigma_{1}(p)\geq\sigma_{v,\varepsilon}(p,x)\geq\sigma_{2}(p)$ for every $p\in\mathbb{R}^{d}$ and every $x\in\Omega$.
For $\varepsilon\in(0,1)$, consider
$\sigma_{v,\varepsilon}(\varepsilon+|Du|)\left[\varepsilon u+F(D^{2}u)\right]=f\text{ in }\Omega.$ (10)
Inspired by ideas in [25], we start by establishing a comparison principle for sub and super-solutions to (10).
Proposition 4 (Comparison Principle).
Let $u\in{\rm USC}(\overline{\Omega})$ be a viscosity sub-solution to (10), and $w\in{\rm LSC}(\overline{\Omega})$ be a viscosity super-solution to (10). Suppose $\Omega$ is a bounded domain, $F$ is degenerate elliptic, and $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$. If $u\leq w$ on $\partial\Omega$, then, $u\leq w$ in $\Omega$.
Proof.
As usual, the proof follows from a contradiction argument. Suppose the statement of the proposition is false. Then there is $x_{0}$ such that
$(u-w)(x_{0})=\max\limits_{x\in\Omega}u-w>0.$
Set $\tau:=(u-w)(x_{0})$. For $\delta>0$, define $\Phi:\overline{\Omega}\times\overline{\Omega}\to\mathbb{R}$ as
$\Phi_{\delta}(x,y):=u(x)-w(y)-\frac{|x-y|^{2}}{2\delta}.$
Denote with $(x_{\delta},y_{\delta})$ the minimiser of $\Phi_{\delta}(x,y)$. Because
$\lim\limits_{\delta\rightarrow 0}\frac{|x_{\delta}-y_{\delta}|^{2}}{2\delta}=0,$
both $x_{\delta}$ and $y_{\delta}$ are in $\Omega$, provided $\delta>0$ is taken sufficiently small; see [22, Lemma 3.1]. From Proposition 1, there are $X,Y$ such that
$\left(\frac{x_{\delta}-y_{\delta}}{\delta},X\right)\in\overline{J}_{2,+}u(x_{\delta})\quad\text{and}\quad\left(\frac{x_{\delta}-y_{\delta}}{\delta},Y\right)\in\overline{J}_{2,-}w(y_{\delta}),$
with
$-\frac{3}{\delta}\begin{pmatrix}I&0\\ 0&I\end{pmatrix}\leq\begin{pmatrix}X&0\\ 0&-Y\end{pmatrix}\leq\frac{3}{\delta}\begin{pmatrix}I&-I\\ -I&I\end{pmatrix}.$
In particular, $X\leq Y$. Hence the degenerate ellipticity of $F$ yields
$\begin{split}\varepsilon\tau^{2}&\leq\varepsilon(u(x_{\delta})-w(y_{\delta}))\\ &\leq\frac{f(x_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}-\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)}.\end{split}$
Combining Remark 1 and Assumption 3 we get
$\min\left(\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right),\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)\right)\geq{\sigma_{2}}(\varepsilon).$
Hence,
$\begin{split}\frac{f(x_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}&-\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)}\leq\frac{f(x_{\delta})-f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}\\ &+\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}-\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)}\end{split}$
Let $\omega_{1}$ and $\omega_{2}$ be the moduli of continuity of $f$ and $\sigma_{v,\varepsilon}$ respectively. Then
$\frac{f(x_{\delta})-f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}+\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}-\frac{f(y_{\delta})}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)}$
$\leq\omega_{1}(|x_{\delta}-y_{\delta}|)\sigma_{2}(\varepsilon)+\left|f\right|_{L^{\infty}(\Omega)}\frac{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)-\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)}{\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},x_{\delta}\right)\cdot\sigma_{v,\varepsilon}\left(\varepsilon+\frac{|x_{\delta}-y_{\delta}|}{\delta},y_{\delta}\right)}$
$\leq\omega_{1}(|x_{\delta}-y_{\delta}|)\sigma_{2}(\varepsilon)+\left\|f\right\|_{L^{\infty}(\Omega)}\frac{\omega_{2}(|x_{\delta}-y_{\delta}|)}{\sigma_{2}(\varepsilon)^{2}}$
Because
$\omega_{1}(|x_{\delta}-y_{\delta}|)\sigma_{2}(\varepsilon)+\left\|f\right\|_{L^{\infty}(\Omega)}\frac{\omega_{2}(|x_{\delta}-y_{\delta}|)}{\sigma_{2}(\varepsilon)^{2}}\longrightarrow 0$
as we take the limit $\delta\to 0$, we conclude $\varepsilon\tau^{2}\leq 0$, which yields a contradiction and completes the proof. The result is then obtained by letting $\delta$ go to zero. ∎
Once the comparison principle is available, we proceed by constructing explicit sub and super-solutions for (10). This is the subject of the next proposition.
Proposition 5 (Existence of global sub and super-solutions).
Let $\Omega$ be a domain satisfying u uniform exterior sphere condition. Suppose $F$ is a degenerate elliptic operator, $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$ and $g\in C(\partial\Omega)$. Suppose further that Assumption 3 holds. Then there is a pair of functions $\overline{w},\underline{w}$ in $C(\overline{\Omega})$ such that, for every $\varepsilon>0$ and $v\in C(\overline{\Omega})$, $\overline{w}$, $\underline{w}$ are viscosity super and sub-solutions to (10) respectively.
Proof.
We will construct only a super-solution to (10) since one analogously obtains a sub-solution. Our argument follows exactly along the same lines as in [25]. In the sequel, we only mention the adjustments necessary to the present setting.
Define functions $w_{1}$ and $w_{y,\eta}$ as in [25, Lemma 2]. From the definition of $w_{1}$, we conclude
$\sigma_{v,\varepsilon}(\varepsilon+|Dw_{1}|,x)[\varepsilon w_{1}F(D^{2}w_{1}(x))]\geq f(x),$ (11)
for every $x\in\Omega$. Also, the definition of $w_{y,\eta}$ ensures that
$\sigma_{v,\varepsilon}(\varepsilon+|Dw_{y,\eta}|)[\varepsilon w_{y,\eta}(x)+F(D^{2}w_{y,\eta}(x))$ $\geq f(x),$
for every $10<\varepsilon,\eta<$ 1, and $x\in\Omega$. Therefore, the functions $w_{y,\eta}$ are super-solutions of (10) for every $10<\varepsilon,\eta<$ 1 and $y\in\partial\Omega$. Thus, the functions
$\widetilde{w}_{y,\eta}(x):=\min(w_{y,\eta}(x),w_{1}(x))$ (12)
are viscosity super-solutions of (10), since the minimum of viscosity super-solutions is a super-solution. In the same manner, we have at last that the function
[…]
is a viscosity super-solution satisfying $\overline{w}=g$ on $\partial\Omega$. ∎
Corollary 1.
Let $\Omega$ be a domain satisfying a uniform exterior sphere condition. Suppose $F$ is degenerate elliptic, $f\in L^{\infty}(\Omega)\cap C(\overline{\Omega})$ and $g\in C(\partial\Omega)$. Furthermore, assume Assumption 3 holds. Then for every $\varepsilon>0$ and $v\in C(\overline{\Omega})$, there is a unique viscosity solution $u_{v,\varepsilon}$ in $C(\overline{\Omega})$, such that $\underline{w}\leq u_{v,\varepsilon}\leq\overline{w}\in\Omega$. There is also $\beta(d,\lambda,\Lambda)>0$, such that for every $\Omega^{\prime}\subset\Omega$,
$\left\|u_{v,\varepsilon}\right\|_{C^{\beta}(\Omega^{\prime})}\leq C,$ (14)
for some $C=(d,\lambda,\Lambda,\left\|u_{v,\varepsilon}\right\|_{L^{\infty}(\Omega)},\left\|f\right\|_{L^{\infty}(\Omega)},{\rm dist}(\Omega^{\prime},\partial\Omega))$
Proof.
To prove the existence of $u_{v,\varepsilon}$, one combines the comparison principle (Proposition 4) and the existence of global sub and super-solutions (Proposition 5) and resort to Perron’s method. To obtain (14) notice that $u_{v,\varepsilon}$ is a viscosity sub-solution of
$\mathcal{M}_{\lambda,\Lambda}^{-}(D^{2}u_{v,\varepsilon})=\left\|f\right\|_{L^{\infty}(\Omega)}+\left\|u_{v,\varepsilon}\right\|_{L^{\infty}(\Omega)}\text{ in }\{|Du_{v,\varepsilon}|>1\}$
and a viscosity super-solution to
$\mathcal{M}_{\lambda,\Lambda}^{+}(D^{2}u_{v,\varepsilon})=-\left\|f\right\|_{L^{\infty}(\Omega)}-\left\|u_{v,\varepsilon}\right\|_{L^{\infty}(\Omega)}\text{in }\{|Du_{v,\varepsilon}|>1\}.$ (15)
A straightforward application of Proposition 2 completes the proof. ∎
Once we have established the existence of a unique viscosity solution $u_{v,\varepsilon}$ to (10) agreeing with $g$ on $\partial\Omega$, we address the parameters $v\in C(\overline{\Omega})$ and $\varepsilon>0$. To handle the functional parameter $v$, we resort to the Schauder fixed point theorem. Define $B\subset C(\overline{\Omega})$ as
$B:=\left\{w\in C(\overline{\Omega})\,|\,\underline{w}\leq w\leq\overline{w}\right\},$ (16)
where $\underline{w}$ and $\overline{w}$ are the sub and super-solutions constructed in Proposition 5. Define also the map $T:B\to C(\overline{\Omega})$ as $Tv:=u_{v,\varepsilon}$. We prove there exists a fixed point $v^{*}\in B$ for $T$.
Proposition 6 (Properties of the map T).
Let $B\subset C(\overline{\Omega})$ be defined as in (16). Define $T:B\rightarrow C(\overline{\Omega})$ as $Tv:=u_{v,\varepsilon}$. Then there exists $v^{\ast}\in B$ such that $Tv^{\ast}=v^{\ast}$.
The proof of Proposition 6 follows along the same lines as in [25, Lemma 3] and is omitted. In what follows, we detail the proof of Theorem 1.
Proof of Theorem 1.
Proposition 6 is tantamount to the existence of $u_{\varepsilon}\in B$ such that
$\sigma_{u_{\varepsilon},\varepsilon}(\varepsilon+|Du_{\varepsilon}|,x)[\varepsilon u_{\varepsilon}+F(D^{2}u_{\varepsilon})]=f\hskip 14.45377pt\mbox{in}\hskip 14.45377pt\Omega,$ (17)
with $u_{\varepsilon}=g$ on $\partial\Omega$.
Corollary 1 builds upon the Arzelà-Ascoli theorem to produce a subsequence $(u_{\varepsilon_{n}})_{n\in\mathbb{N}}$, with $\varepsilon_{n}<1/n$, and a function $u\in B$ such that $u_{\varepsilon_{n}}\rightarrow u$ in $C(\Omega)$, as $n\to\infty$. Since $\sigma_{{v,\varepsilon}}$ converges to $\sigma_{1}\chi_{\{u>0\}}+\sigma_{2}\chi_{\{u<0\}}$ uniformly on compact subsets of ($\left(\{u>0\}\cup\{u<0\}\right)\cap\Omega$, a standard argument on the stability of viscosity solutions allows us to conclude that $u$ solves (1) in ($\left(\{u>0\}\cup\{u<0\}\right)\cap\Omega$. Moreover, since $\sigma_{1}\geq\sigma_{v,\varepsilon}\geq\sigma_{2}$, $u$ is also a viscosity sub-solution to (7) and a viscosity super-solution to (8) in $\Omega$. ∎
Remark 2 (Existence of solutions for the pure equation).
We notice the arguments above can be easily adjusted to provide an existence result for the problem
[…]
provided $\Omega\subset\mathbb{R}^{d}$ satisfies a uniform exterior sphere condition and $g\in C(\partial\Omega)$. The main modification is in the choice of the approximated problem. Indeed, we should consider
[…]
and establish a comparison principle along the same lines as before. Then building sub and super-solutions also following the previous rationale, would allow one to evoke Perron’s method and conclude the argument.
Remark 3 (Existence of globally Lipschitz-continuous solutions).
We believe it is possible to change the sub and super-solutions in Proposition 5 according to the model proposed in [12, Lemma 2.2]. Under such a construction, and resorting to the conclusion of Theorem 2, we conclude the existence of a globally Lipschitz-continuous viscosity solution to (1).
4 Local $C^{1}$-regularity estimates
@@
$\min\left(\sigma_{1}(|q+Du|)F(D^{2}u),\sigma_{2}(|q+Du|)F(D^{2}u)\right)\leq C_{0}\hskip 14.45377pt\mbox{in}\hskip 14.45377ptB_{1}$ (19)
@@
$\max\left(\sigma_{1}(|q+Du|)F(D^{2}u),\sigma_{2}(|q+Du|)F(D^{2}u)\right)\geq-C_{0}\hskip 14.45377pt\mbox{in}\hskip 14.45377ptB_{1},$ (20)
where $q\in\mathbb{R}^{d}$ is fixed, though arbitrary. Suppose Assumptions 2 and 3 are in force. Suppose further $|q|\leq A_{0}$, for some constant $A_{0}>1$. Then there exists $\beta\in(0,1)$ such that $u\in C^{\beta}_{\rm loc}(B_{1})$. In addition, for every $\tau\in(0,1)$ there exists $C_{\tau}>0$ for which
@@
$\overline{\sigma}:=\min\left(\sigma_{1}(A_{0}),\sigma_{2}(A_{0})\right)\leq\min\left(\sigma_{1}(|q+p|),\sigma_{2}(|q+p|)\right).$
@@
By setting $\gamma:=2A_{0}$ in Proposition 2 the result follows. ∎
We continue with an application of the maximum principle to produce Hölder continuity for a solution to (19)-(20) in case $|q|\geq A_{0}$. Our argument follows along the same lines as in [25, Proposition 5].
@@
Let $u\in C(B_{1})$ be a viscosity solution to (19)-(20). Suppose Assumptions 2 and 3 are in force. Then $u\in C^{\beta}_{\rm loc}(B_{1})$, for some universal constant $\beta\in(0,1)$. In addition, for every $\rho\in(0,1)$ there exists $C>0$ such that
@@
We argue as in the proof of [25, Proposition 5]; in the sequel, we omit most of the details, stressing the main differences with respect to the argument in that paper. We split the proof into four steps.
Step 1 - Fix $20<r<(1-\tau)/$ 2 and consider the modulus of continuity $\omega(t):=t-t^{2}/2$. Define the quantity
$L:=\sup_{x,y\in B_{r}(x_{0})}\left(u(x)-u(y)-L_{1}\omega(|x-y|)-L_{2}\left(|x-x_{0}|^{2}+|y-x_{0}|^{2}\right)\right).$
As usual, our goal is to choose $L_{1}$ and $L_{2}$ such that $L\leq 0$ for every $x_{0}\in B_{\tau}$. Suppose such a choice is not possible. It is tantamount to say there exists $x_{0}$ such that $L>0$ regardless of the choice of $L_{1}$ and $L_{2}$. Arguing as in [25, Proposition 5], we find points $(X,p_{x},x),\,(Y,p_{y},y)\in S(d)\times\mathbb{R}^{d}\times B_{r}(x_{0})$, and a constant $\iota>0$, such that
$\mathcal{M}_{\lambda,\Lambda}^{-}(X-Y)\geq 4\lambda L_{1}-(\lambda+(d-1)\Lambda)(4L_{2}+2\iota),$ (21)
$\min\left(\sigma_{1}(|q+p_{x}|)F(X),\sigma_{2}(|q+p_{x}|)F(X)\right)\leq C_{0},$ (22)
$\max\left(\sigma_{1}(|q+p_{y}|)F(Y),\sigma_{2}(|q+p_{y}|)F(Y)\right)\leq C_{0},$ (23)
@@
$F(X)\geq F(Y)+\mathcal{M}_{\lambda,\Lambda}^{-}(X-Y).$ (24)
@@
$4\lambda L_{1}\leq\left(\lambda+(d-1)\Lambda\right)\left(4L_{2}+2\iota\right)+C_{0}\left(\frac{1}{\sigma_{i}(|q+p_{j}|)}+\frac{1}{\sigma_{k}(|q+p_{\ell}|)}\right),$ (25)
@@
where $L_{2}:=(4\sqrt{2}/r)^{2}$, we conclude there exists $a>0$ such that
@@
Let $u\in C(B_{1})$ be a viscosity solution to (19)-(20). Suppose Assumptions 2 and 3 are in force. Let $\alpha_{0}\in(0,1)$ be the exponent for the Krylov-Safonov regularity theory available for $F=0$. For every $\delta>0$ there exists $\varepsilon>0$ such that, if $C_{0}\leq\varepsilon$ then one can find $h\in C^{1,\alpha_{0}}_{\rm loc}(B_{9/10})$ satisfying
@@
$p(x):=u_{\infty}(y)+{\rm b}\cdot(x-y)+\frac{1}{2}(x-y)^{T}M(x-y).$
@@
$\min\left(\sigma_{1}^{n}\left(\left|{\rm b}+q_{n}\right|\right)F_{n}(M),\sigma_{2}^{n}\left(\left|{\rm b}+q_{n}\right|\right)F_{n}(M)\right)\leq C_{n}.$ (26)
@@
$\varphi(x):=\kappa\sup_{e\in\mathbb{S}^{d-1}}\left\langle P_{E}x,e\right\rangle+\frac{1}{2}x^{T}Mx,$
@@
$\varphi_{e}(x):=\kappa\left\langle P_{E}x,e\right\rangle+\frac{1}{2}x^{T}Mx.$
@@
$\min\left(\sigma_{1}^{n}\left(\left|\kappa e+Mx_{n}+q_{n}\right|\right)F_{n}(M),\sigma_{2}^{n}\left(\left|\kappa e+Mx_{n}+q_{n}\right|\right)F_{n}(M)\right)\leq C_{n}.$ (27)
@@
$e:=\frac{Mx_{n}}{\left|Mx_{n}\right|}$
@@
$F_{n}(M)\leq\frac{C_{n}}{\sigma_{1}^{n}\left(\kappa/2\right)}\qquad\mbox{or}\qquad F_{n}(M)\leq\frac{C_{n}}{\sigma_{2}^{n}\left(\kappa/2\right)},$ (28)
@@
Step 6 - Because $F_{\infty}(D^{2}u_{\infty})=0$ in $B_{1}$, standard regularity results imply that $u_{\infty}\in C^{1,\alpha_{0}}_{loc}(B_{1})$ with estimates, where $\alpha_{0}\in(0,1)$ is the (universal) exponent stemming from the Krylov-Safonov theory. By taking $h:=u_{\infty}$ one finds a contradiction and concludes the proof. ∎
We use Proposition 9 to construct a sequence of approximating hyperplanes. The difference between the solution and such approximating hyperplanes behave in a precise geometric fashion. Such a geometric control of the difference between the solution and a hyperplane is key to differentiability.
4.1 Proof of Theorem 2
We start by introducing $\gamma(t):=t\sigma_{2}(t)$; because $t\mapsto t\sigma_{2}(t)$ is a bijective map, it has an inverse. Set $\omega(t):=\gamma^{-1}(t)$. We use $\omega(\cdot)$ to examine an alternative.
For $\alpha_{0}\in(0,1)$ as in Proposition 9, suppose first
$\frac{t^{\alpha_{0}}}{\omega(t)}\longrightarrow 0,\qquad\mbox{as}\quad t\to 0,$
we choose $r\in(0,1)$ and define $\mu_{1}>0$ such that
$2Cr^{\alpha_{0}}=\omega(r)=:\mu_{1},$ (29)
where $C>0$ is the universal constant in Proposition 9. On the contrary, suppose that
$\frac{t^{\alpha_{0}}}{\omega(t)}\longrightarrow M,\qquad\mbox{as}\quad t\to 0,$
where $M\in\mathbb{R}$ is some positive constant. In this case, set
$\mu_{1}:=r^{\frac{\alpha_{0}}{2}}=\frac{1}{2C},$ (30)
where $\alpha_{0}\in(0,1)$ and $C>0$ are once again the constants in Proposition 9. We notice both (29) and (30) imply
$2Cr^{1+\alpha_{0}}=\mu_{1}r$ (31)
We proceed by defining the constant $\theta=\frac{r}{\mu_{1}}$ and considering the sequence
$(a_{k})_{k\in\mathbb{N}}:=\left(\sigma_{2}^{-1}(\theta^{k})\right)_{k\in\mathbb{N}}.$
Under Assumption 4, the inverse $\sigma_{2}^{-1}$ is Dini-continuous; as a consequence, the sequence $\left(\sigma_{2}^{-1}(\theta^{k})\right)_{k\in\mathbb{N}}$ is summable, and $(a_{k})_{k\in\mathbb{N}}\in\ell^{1}$. Now, we resort to Proposition 3. For $0<\delta<\frac{1}{4}$, we set $10<\varepsilon<$ 1 as
$\varepsilon:=\frac{1}{1+\delta}.$

### v2

@@ 2. Preliminaries
3. Local $C^{1}$-regularity estimates
@@
(7) $\min\left(\sigma_{1}(|q+Du|)F(D^{2}u),\sigma_{2}(|q+Du|)F(D^{2}u)\right)\leq C_{0}\hskip 14.45377pt\mbox{in}\hskip 14.45377ptB_{1}$
@@
(8) $\max\left(\sigma_{1}(|q+Du|)F(D^{2}u),\sigma_{2}(|q+Du|)F(D^{2}u)\right)\geq-C_{0}\hskip 14.45377pt\mbox{in}\hskip 14.45377ptB_{1},$
where $q\in\mathbb{R}^{d}$ is fixed, though arbitrary. Suppose Assumptions 1 and 2 are in force. Suppose further $|q|\leq A_{0}$, for some constant $A_{0}>1$. Then there exists $\beta\in(0,1)$ such that $u\in C^{\beta}_{\rm loc}(B_{1})$. In addition, for every $\tau\in(0,1)$ there exists $C_{\tau}>0$ for which
@@
$\overline{\sigma}\coloneqq\min\left(\sigma_{1}(A_{0}),\sigma_{2}(A_{0})\right)\leq\min\left(\sigma_{1}(|q+p|),\sigma_{2}(|q+p|)\right).$
@@
By setting $\gamma\coloneqq 2A_{0}$ in Proposition 2 the result follows. ∎
We continue with an application of the maximum principle to produce Hölder continuity for a solution to (7)-(8) in case $|q|\geq A_{0}$. Our argument follows along the same lines as in [11, Proposition 5].
@@
Let $u\in C(B_{1})$ be a viscosity solution to (7)-(8). Suppose Assumptions 1 and 2 are in force. Then $u\in C^{\beta}_{\rm loc}(B_{1})$, for some universal constant $\beta\in(0,1)$. In addition, for every $\rho\in(0,1)$ there exists $C>0$ such that
@@
We argue as in the proof of [11, Proposition 5]; in the sequel, we omit most of the details, stressing the main differences with respect to the argument in that paper. We split the proof into four steps.
Step 1 - Fix $20<r<(1-\tau)/$ 2 and consider the modulus of continuity $\omega(t)\coloneqq t-t^{2}/2$. Define the quantity
$L\coloneqq\sup_{x,y\in B_{r}(x_{0})}\left(u(x)-u(y)-L_{1}\omega(|x-y|)-L_{2}\left(|x-x_{0}|^{2}+|y-x_{0}|^{2}\right)\right).$
As usual, our goal is to choose $L_{1}$ and $L_{2}$ such that $L\leq 0$ for every $x_{0}\in B_{\tau}$. Suppose such a choice is not possible. It is tantamount to say there exists $x_{0}$ such that $L>0$ regardless of the choice of $L_{1}$ and $L_{2}$. Arguing as in [11, Proposition 5], we find points $(X,p_{x},x),\,(Y,p_{y},y)\in S(d)\times\mathbb{R}^{d}\times B_{r}(x_{0})$, and a constant $\iota>0$, such that
(9) $\mathcal{M}_{\lambda,\Lambda}^{-}(X-Y)\geq 4\lambda L_{1}-(\lambda+(d-1)\Lambda)(4L_{2}+2\iota),$
(10) $\min\left(\sigma_{1}(|q+p_{x}|)F(X),\sigma_{2}(|q+p_{x}|)F(X)\right)\leq C_{0},$
(11) $\max\left(\sigma_{1}(|q+p_{y}|)F(Y),\sigma_{2}(|q+p_{y}|)F(Y)\right)\leq C_{0},$
@@
(12) $F(X)\geq F(Y)+\mathcal{M}_{\lambda,\Lambda}^{-}(X-Y).$
@@
(13) $4\lambda L_{1}\leq\left(\lambda+(d-1)\Lambda\right)\left(4L_{2}+2\iota\right)+C_{0}\left(\frac{1}{\sigma_{i}(|q+p_{j}|)}+\frac{1}{\sigma_{k}(|q+p_{\ell}|)}\right),$
@@
where $L_{2}\coloneqq(4\sqrt{2}/r)^{2}$, we conclude there exists $a>0$ such that
@@
Let $u\in C(B_{1})$ be a viscosity solution to (7)-(8). Suppose Assumptions 1 and 2 are in force. Let $\alpha_{0}\in(0,1)$ be the exponent for the Krylov-Safonov regularity theory available for $F=0$. For every $\delta>0$ there exists $\varepsilon>0$ such that, if $C_{0}\leq\varepsilon$ then one can find $h\in C^{1,\alpha_{0}}_{\rm loc}(B_{9/10})$ satisfying
@@
$p(x)\coloneqq u_{\infty}(y)+{\rm b}\cdot(x-y)+\frac{1}{2}(x-y)^{T}M(x-y).$
@@
(14) $\min\left(\sigma_{1}^{n}\left(\left|{\rm b}+q_{n}\right|\right)F_{n}(M),\sigma_{2}^{n}\left(\left|{\rm b}+q_{n}\right|\right)F_{n}(M)\right)\leq C_{n}.$
@@
$\varphi(x)\coloneqq\kappa\sup_{e\in\mathbb{S}^{d-1}}\left\langle P_{E}x,e\right\rangle+\frac{1}{2}x^{T}Mx,$
@@
$\varphi_{e}(x)\coloneqq\kappa\left\langle P_{E}x,e\right\rangle+\frac{1}{2}x^{T}Mx.$
@@
(15) $\min\left(\sigma_{1}^{n}\left(\left|\kappa e+Mx_{n}+q_{n}\right|\right)F_{n}(M),\sigma_{2}^{n}\left(\left|\kappa e+Mx_{n}+q_{n}\right|\right)F_{n}(M)\right)\leq C_{n}.$
@@
$e\coloneqq\frac{Mx_{n}}{\left|Mx_{n}\right|}$
@@
(16) $F_{n}(M)\leq\frac{C_{n}}{\sigma_{1}^{n}\left(\kappa/2\right)}\qquad\mbox{or}\qquad F_{n}(M)\leq\frac{C_{n}}{\sigma_{2}^{n}\left(\kappa/2\right)},$
@@
Step 6 - Because $F_{\infty}(D^{2}u_{\infty})=0$ in $B_{1}$, standard regularity results imply that $u_{\infty}\in C^{1,\alpha_{0}}_{loc}(B_{1})$ with estimates, where $\alpha_{0}\in(0,1)$ is the (universal) exponent stemming from the Krylov-Safonov theory. By taking $h\coloneqq u_{\infty}$, one finds a contradiction and concludes the proof. ∎
We use Proposition 6 to construct a sequence of approximating hyperplanes. The difference between the solution and such approximating hyperplanes behaves in a precise geometric fashion. Such a geometric control of the difference between the solution and a hyperplane is key to differentiability.
3.1. Proof of Theorem 1
For $\alpha_{0}\in(0,1)$ as in Proposition 6, choose $r\in(0,1)$ and $\mu_{1}>0$ such that
(17) $2Cr^{1+\alpha_{0}}=\mu_{1}r,$
where $C>0$ is the universal constant in Proposition 6.
We proceed by defining the constant $\theta=\frac{r}{\mu_{1}}$ and considering the sequence $(a_{k})_{k\in\mathbb{N}}$ defined as
$a_{k}\coloneqq\max\left(\sigma_{1}^{-1}(\theta^{k}),\sigma_{2}^{-1}(\theta^{k})\right).$
Under Assumption 2, we conclude $(a_{k})_{k\in\mathbb{N}}\in\ell^{1}$. Now, we resort to Proposition 3. For $0<\delta<\frac{1}{4}$, we set $10<\varepsilon<$ 1 as
$\varepsilon\coloneqq\frac{1}{1+\delta}.$